OpenStudy (anonymous):

The area of a rectangle is 36 sq cm. The lengths of the sides are whole numbers of centimenters. WHat is the greatest perimter the rectangle can have , in cm?

OpenStudy (amistre64):

a square

OpenStudy (amistre64):

2(1+36) = 74 2(2+18) = 40 2(3+12) = 30 2(4+9) = 26 2(6+6) = 24 ..... hmm i mighta got my thoughts mixed up :)

OpenStudy (anonymous):

Please Explain, as I don't fully understand your thoughts.

OpenStudy (anonymous):

yep - not a square bu a long thin rectangle

OpenStudy (amistre64):

the greatest are is a square :) ... so many memories, not enough place to put them lol

OpenStudy (amistre64):

Area = s1 * s2 Perimeter = 2(s1) + 2(s2) = 2(s1 + s2)